OpenStudy (anonymous):

Factor in ( + ) ( - ) form 32v^2-2

OpenStudy (anonymous):

\[32v^2 -2 = 2(16v^2 -1)\] @OliviannaC

jhonyy9 (jhonyy9):

@Adjax yes but you can continue it again so inside parentheses

jhonyy9 (jhonyy9):

2(16v^2 -1) = 2(4v -1)(4v+1)